Solid A and solid B are similar. The surface area of solid A is 675 m2 and the surface area of solid B is 432 m2. If the volume of solid B is 960 m3, find the volume of solid A.


1
Expert's answer
2021-03-04T01:31:49-0500


Vb = the volume of solid B is =960m3

Sa = the surface area of A =675 m2

Sb = the surface area of B=432 m2

volume

= Area * Hight (h)

Sb * h = 960 m 3

h=

k^2= Sa / Sb =675/432=1.56

k=1.25

h’=hk

h’=2.22*1.25

h’=2.775

Va= Sa*h’=675*2.775

Sa=1873.125 m3